OpenStudy (anonymous):

Convert the complex number into its polar representation: -4i

OpenStudy (maheshmeghwal9):

\[complex no. is -4i=(0-4i)\]

OpenStudy (maheshmeghwal9):

polar representation is \[re^{i \theta}=r(\cos \theta+isin \theta)=rcos \theta+ri \sin \theta=x+iy=a complex no.\] put the values u can do it

OpenStudy (maheshmeghwal9):

where \[r=\sqrt{x^2+y^2}\]

OpenStudy (maheshmeghwal9):

& \[\theta =\tan^{-1} (y/x)\]

Directrix (directrix):